OpenStudy (anonymous):

1.for which values of x and y is line p parallel to line q? (x = 5, y =3) (x = 1, y = 5) (x =3, y=5)(x=3,y=6) 2. Which of the following equations would graph a line parallel to 3y = 2x - x +5? (3y=x+1)(y=x+1)(y=2/3x+1)(3/2y=x+1) 3. Plans for a stadium are drawn on a coordinate grid. One wall lies on the line y = 2x + 3. A perpendicular wall passes through the point (4,-7). Write the equation(y=-1/2x-5)(y=2x^2-15)(y=1/2x-9)(y=^2-2x+1) 4. Write an equation in point-slop form of the line through point J(4,1) with slope -4.

OpenStudy (anonymous):

For 1

OpenStudy (anonymous):

I tried my best to do some of the problems but i fell short my best guesses are 1. D 2. B 3. A 4. idk

OpenStudy (anonymous):

1. Line p line q ? 2. 3y = x + 5 (1(--> y = x/3 + 5. Line 3y = x + 1 would parallel to line (1) 3. Wall line: y = 2x + 3 (1). Any line perpendicular to (1) will have slope -1/2. y = -x/2 + b. Write that this line passes at point (4, -7). -7 = -(4)/2 + b --> b = -7 + 2 = -5 Equation y = -x/2 -5 4. y - -4x + b Find b by writing that this line passes at point (4, 1): 1 = -4 + b.